# Conformalized Interval Arithmetic with Symmetric Calibration

### v1

@@ 6.1. Multiple Test
Conformal prediction has been applied to multiple testing [10, 13, 5, 4, 2]. While our present work also deal with multiple random variables, CIA provides only one confidence interval for the sum of these random variables, so our method is essentially different from these multiple test-based approaches.
6.2. Half Sampling
@@
In Theorem 4.1, the conclusion can be restated as follows: the $(1-\alpha)$-th quantile of the score function values in the calibration set is larger than a proportion of $1-\alpha$ of the score function values in the test set. At a higher level, this is a comparison between the $(1-\alpha)$-th quantile of half of the samples and the entire sample set. In this regard, symmetric calibration can be linked to the balanced half-sample method [15, 32], which was popular from 1970 to 2000. The balanced half-sample method shares the same idea as symmetric calibration, as it compares the statistic from the half-sample with the entire sample. It is possible that further theory about symmetric calibration can be developed by building upon classical half-sample theory.
6.3. Fair Conformal Prediction
@@
Fair conformal prediction [18, 17, 36] and group-conditional conformal prediction [28, 9, 25] aim to equalize coverage across various groups characterized by group attributes or auxiliary data. This idea is related to the stratified conformal prediction discussed in this paper. The procedures in these methods are very similar. However, the motivation behind stratified conformal prediction is to improve the efficiency of the conformal prediction, while group-conditional conformal prediction is designed to ensure fairness among different groups.
6.4. Robust Route Planning
@@
As mentioned in the introduction, our method can be applied to path cost prediction. We will further discuss this application in Section 7.2. In [34, 26], the authors introduce Conformal-Predict-Then-Optimize (CPO) for linear programming by predicting edge weights and minimizing upper bounds of path costs for optimal robust solutions. While our method can generate conformal predictions for algorithm-selected paths, choosing the path with the smallest upper bound depends on the score function, compromising the guarantee of $1-\alpha$ coverage. Additionally, applying CPO to our problem is inefficient, as it only provides conformal predictions for the joint distribution of edge weights. We are also aware of recent works [6, 14, 33] which address decision making under uncertainty.
7. Application and Experiment
@@
In this section, we will introduce two main applications of the proposed methods. Then we will analyze the empirical performance of our algorithms on public datasets. The code is publicly available at GitHub.
7.1. Application 1: Group Average Prediction
@@ 7.1.1. Datasets
    Bike Sharing [8]: This dataset is used to investigate the factors influencing bike rental demand. The grouping features are season, workingday, and weather, which have 272 unique value combinations out of 10886 samples.
  2. (2)
@@
    Community Crime [27]: This dataset is used to predict the per capita violent crime rate of a community using its demographic features. The grouping features are state and county, which have 350 unique value combinations out of 1994 samples.
  3. (3)
@@
    Medical Expenditure Panel Survey [1]: This dataset is used to predict the utilization of medical services based on various features. The grouping features are age, race, and marriage, which have 302, 302, and 301 unique value combinations out of 15785, 17541, and 15656 samples for the years 2019, 2020, and 2021, respectively.
@@
Following similar procedures in [30], we standardize the response variables $Y$ for all datasets. We use $70\%$ of data for training the quantile regression model and $30\%$ for calibration and testing.
7.2. Application 2: Path Cost Prediction
@@
As discussed in the introduction, our method has a valuable application in route planning problems. We consider a transductive setting in edge regression problems, where $\mathcal{I}$ represents the set of edge indices. In the experiment, we assume that there are edges with unknown labels, and we denote the set of these edges as $\mathcal{I}_{\textrm{test}}$. Similar to Algorithm 1, we can hold out a set of edges with known labels as the calibration set $\mathcal{I}_{\textrm{cal}}$. Utilizing node and/or edge features, the network structure, and the observed edge labels (costs), we train a Graph Neural Network (GNN) to predict the labels of the calibration set and test set. The transductive setting of GNN is introduced in [11], where they consider the prediction of node labels. However, more recent work [19] which focuses on edge prediction is more suitable for this particular application set.
@@
While the choice of $S_{k}$’s can be quite general, the most interesting case in the traffic network data is when we use the edges on the shortest path as the set of edges $S_{k}$. First, we randomly sample two nodes $(s,t)$ as the source and target nodes, respectively. Then, we use $y_{i}$ for $i\in\mathcal{I}_{\textrm{train}}$ and $\widehat{y}_{i}$ for $i\in\mathcal{I}_{\textrm{cal}}\cup\mathcal{I}_{\textrm{test}}$ as the cost of each edge. Next, we find the shortest path using Dijkstra’s algorithm. By sampling $K+1$ pairs of $(s,t)$, the indices of edges on the path form an index set $S_{(s,t)}$. Given the index sets and the predictions $\widehat{y}_{i}$ for calibration samples and test samples, we can apply Algorithm 1 or Algorithm 2 to obtain the conformal prediction.
@@
In this setting, the index sets can overlap, meaning that different paths can contain the same edge. However, in a large network without bottlenecks, two shortest paths with random sources and targets share the same edge with a small probability. Therefore, Theorem 5.2 can guarantee that the coverage of the prediction set is very close to $1-\alpha$.
@@
Road Traffic in Anaheim and Chicago [3]. This dataset is used to predict the traffic flow along certain roads based on the traffic flow along other roads. The predicted traffic flow is then utilized as edge cost for shortest path planning. This scenario corresponds to the situation of subsets with overlaps, where the edges (roads) in different subsets (shortest paths) may overlap with each other. We collect 2000 shortest paths by randomly sampling $(s,t)$ pairs. The Anaheim dataset consists of 413 nodes and 858 edges, and the Chicago dataset consists of 541 nodes and 2150 edges. We adopt a similar procedure from [12, 11], and allocate 50%, 10%, and 40% for training, validation, and calibration and testing.
7.3. Baseline
@@
This method corresponds to the approach described in Section 3. To predict $\sum_{i\in S^{\textrm{test}}_{K+1}}y_{i}$, we sample $K$ groups of samples from the calibration sets, where each group contains $|S^{\textrm{test}}_{K+1}|$ samples. Let $S_{1},\dots,S_{K}$ be the index sets of these groups. The prediction set is then given by (1).
It is important to note that in this method, the indices of $S_{1},\dots,S_{K}$ may not belong to the same class, which means that this approach may not be appropriate in all cases. However, the method is always applicable, regardless of the class composition of the sampled groups.
@@
Then, the prediction set for $\sum_{i\in S^{\textrm{test}}_{K+1}}y_{i}$ is given by
$\left[\sum_{i\in S^{\textrm{test}}_{K+1}}\widehat{y}_{i}+z_{\frac{\alpha}{2}}\sqrt{|S^{\textrm{test}}_{K+1}|}\widehat{\sigma},\sum_{i\in S^{\textrm{test}}_{K+1}}\widehat{y}_{i}+z_{1-\frac{\alpha}{2}}\sqrt{|S^{\textrm{test}}_{K+1}|}\widehat{\sigma}\right],$
@@
The Bonferroni correction has been employed in various conformal methods [10, 13, 7]. As mentioned in Section 2.2, it is also applicable to our problem. While the Bonferroni correction is a valid approach, it often lacks efficiency in the context of our problem.
7.4. Results
@@
Figure 2. Results for constructing prediction sets for subsets without overlaps (Section 7.1). CIA achieves guaranteed $1-\alpha$ coverage as well as optimal efficiency on various datasets.
Figure 3. Results for constructing prediction sets for subsets with overlaps (Section 7.2). CIA achieves close to $1-\alpha$ coverage and optimal efficiency on two road traffic datasets.
For these five datasets, the left plots present the desired miscoverage rate $\alpha$ versus the true coverage rate. The closer the curve aligns with the line $1-\alpha$, the easier it is for the method to achieve the desired coverage. Our proposed method, CIA, is very close to the line $1-\alpha$. There is a small gap in the result of the Anaheim dataset. This result aligns with Theorem 5.2. Compared to the Chicago dataset, Anaheim is a smaller traffic network. The probability of two shortest paths overlapping each other can affect the validity. This issue is much less obvious in a larger traffic network like Chicago. Recall that the group-sample conformal method is proposed in Section 3. This is also a conformal prediction method, so its validity is also close to the desired line $1-\alpha$ in the results of Figure 2. However, the validity is much less reliable in the results of Figure 3. This is because edge weights from random samples are not exchangeable with the edge weights on the shortest path. The method of Bonferroni correction has good validity in the example of the community dataset. The reason is that under the current setting of calibration and test set ratio, each subset contains one or two samples, which is very close to the case of classical conformal prediction. In other datasets, Bonferroni correction has too high coverage. For normal confidence intervals, since the assumption is not true in general, the coverage is much lower than the desired one. In conclusion, our proposed method has the most reliable validity in all datasets that we have studied.
@@
8. Conclusion
In this paper, we have introduced conformalized interval arithmetic with symmetric calibration, to tackle the challenges of conformal prediction for interdependent data. Our approach offers methodological contributions and theoretical guarantees for achieving the desired coverage under reasonable exchangeability assumptions. Extensive experiments on real-world datasets demonstrate the validity and efficiency of our method, outperforming others even when they fail to ensure valid coverage.
Future works include using novel conformal score functions [20, 22, 23] and weighted aggregation [24] to improve the efficiency. We hope our core ideas and techniques will foster the development of more robust conformal prediction methods across various domains.
@@ Appendix A Proof of Theorem 1
We first consider a fixed permutation and a fixed partition of calibration and test set. Let $r^{\textrm{cal}}_{K+1}$ be the rank of $s^{\textrm{cal}}_{K+1}$ in $s^{\textrm{cal}}_{k}$ for $k\in[K+1]$. We exchange the calibration set and test set in $S_{K+1}$, i.e., we swap $S^{\textrm{test}}_{K+1}$ and $S^{\textrm{cal}}_{K+1}$. Under the symmetry assumption on the calibration indices and test indices, this swapping procedure corresponds two possible partitions with the same probability. Let $r^{\textrm{test}}_{K+1}$ be the rank of $s^{\textrm{test}}_{K+1}$ in $\{s^{\textrm{cal}}_{k}:k=1,\dots,K\}\cup\{s^{\textrm{test}}_{K+1}\}$ after swapping. Under the assumption of Theorem 1, the scores of the calibration set remains the same. Hence $r^{\textrm{test}}_{K+1}=r^{\textrm{cal}}_{K+1}$. $r^{\textrm{test}}_{K+1}$ has the same distribution as $r^{\textrm{cal}}_{K+1}$. Hence,
@@
The event $r^{\textrm{test}}_{K+1}\leq\lceil(1+K)(1-\alpha)$ is equivalent to $\sum_{i\in S^{\textrm{test}}_{K+1}}y_{i}\in\mathcal{C}_{(}S^{\textrm{test}}_{K+1})$ using the definition of $Q_{1-\alpha}$. This completes the proof.
@@ Appendix B Proof of Theorem 2
Appendix C Analysis of CIA on Overlapping Subsets

### v2

@@ 6 Related Work
Conformal prediction has been applied to multiple testing (Fisch et al. 2021; Jin and Candès 2023; Bates et al. 2023; Bashari et al. 2024; Angelopoulos et al. 2024). While our present work also deal with multiple random variables, CIA provides only one confidence interval for the sum of these random variables, so ours is essentially different from these multiple test-based approaches.
Half Sampling.
@@
In Theorem 1, the conclusion can be restated as follows: the $(1-\alpha)$-th quantile of the score function values in the calibration set is larger than a proportion of $1-\alpha$ of the score function values in the test set. At a higher level, this is a comparison between the $(1-\alpha)$-th quantile of half of the samples and the entire sample set. In this regard, symmetric calibration can be linked to the balanced half-sample method (Kish and Frankel 1970; Shao and Wu 1992), which was popular from 1970 to 2000. The balanced half-sample method shares the same idea as symmetric calibration, as it compares the statistic from the half-sample with the entire sample. It is possible that further theory about symmetric calibration can be developed by building upon classical half-sample theory.
Fair Conformal Prediction.
@@
Fair conformal prediction (Lu et al. 2022; Liu et al. 2022; Wang et al. 2024) and group-conditional conformal prediction (Romano et al. 2020; Feldman, Bates, and Romano 2021; Melki et al. 2023; Plassier et al. 2024) aim to equalize coverage across various groups characterized by group attributes or auxiliary data. This idea is related to the stratified conformal prediction discussed in this paper. The procedures in these methods are very similar. However, the motivation behind stratified conformal prediction is to improve the efficiency of the conformal prediction, while group-conditional conformal prediction is designed to ensure fairness among different groups.
Robust Route Planning.
@@
As mentioned in the introduction, our method can be applied to path cost prediction. We will further discuss this application in Section 7.2. In (Sun, Liu, and Li 2023; Patel, Rayan, and Tewari 2024), the authors introduce Conformal-Predict-Then-Optimize (CPO) for linear programming by predicting edge weights and minimizing upper bounds of path costs for optimal robust solutions. While our method can generate conformal predictions for algorithm-selected paths, choosing the path with the smallest upper bound depends on the score function, compromising the guarantee of $1-\alpha$ coverage. Additionally, applying CPO to our problem is inefficient, as it only provides conformal predictions for the joint distribution of edge weights. We are also aware of recent works (Bertsimas, Gupta, and Kallus 2018; Chassein, Dokka, and Goerigk 2019; Johnstone and Cox 2021; Stanton, Maddox, and Wilson 2023; Sadana et al. 2024) which address decision making under uncertainty and methods which tackle adversary data (Gendler et al. 2021; Bastani et al. 2022; Ghosh et al. 2023) and distribution shifts (Tibshirani et al. 2019; Prinster et al. 2024; Zhao et al. 2024) in conformal prediction.
7 Application and Experiment
@@
In this section, we will introduce two main applications of the proposed methods. Then we will analyze the empirical performance of our algorithms on public datasets.
7.1 Application 1: Group Average Prediction
@@
    Bike Sharing (Fanaee-T 2013): This dataset is used to investigate the factors influencing bike rental demand. The grouping features are season, workingday, and weather, which have 272 unique value combinations out of 10886 samples.
  2. 2.
@@
    Community Crime (Redmond 2009): This dataset is used to predict the per capita violent crime rate of a community using its demographic features. The grouping features are state and county, which have 350 unique value combinations out of 1994 samples.
  3. 3.
@@
    Medical Expenditure Panel Survey (AfHRaQ 2021): This dataset is used to predict the utilization of medical services based on various features. The grouping features are age, race, and marriage, which have 302, 302, and 301 unique value combinations out of 15785, 17541, and 15656 samples for the years 2019, 2020, and 2021, respectively.
@@
Following similar procedures in (Sesia and Candès 2020), we standardize the response variables $Y$ for all datasets. We use $70\%$ of data for training the quantile regression model and $30\%$ for calibration and testing.
7.2 Application 2: Path Cost Prediction
@@
As discussed in the introduction, our method has a valuable application in route planning problems. We consider a transductive setting in edge regression problems, where $\mathcal{I}$ represents the set of edge indices. In the experiment, we assume that there are edges with unknown labels, and we denote the set of these edges as $\mathcal{I}_{\textrm{test}}$. Similar to Algorithm 1, we can hold out a set of edges with known labels as the calibration set $\mathcal{I}_{\textrm{cal}}$. Utilizing node and/or edge features, the network structure, and the observed edge labels (costs), we train a Graph Neural Network (GNN) to predict the labels of the calibration set and test set. The transductive setting of GNN is introduced in (Huang et al. 2024), where they consider the prediction of node labels. However, more recent works (Luo and Colombo 2024; Zhao, Kang, and Cheng 2024) which focus on edge prediction are more suitable for this particular application set.
@@
While the choice of $S_{k}$’s can be quite general, the most interesting case in the traffic network data is when we use the edges on the shortest path as the set of edges $S_{k}$. First, we randomly sample two nodes $(s,t)$ as the source and target nodes, respectively. Then, we use $y_{i}$ for $i\in\mathcal{I}_{\textrm{train}}$ and $\widehat{y}_{i}$ for $i\in\mathcal{I}_{\textrm{cal}}\cup\mathcal{I}_{\textrm{test}}$ as the cost of each edge. Next, we find the shortest path using Dijkstra’s algorithm. By sampling $K$ pairs of $(s,t)$, the indices of edges on the path form an index set $S_{(s,t)}$. Given the index sets and the predictions $\widehat{y}_{i}$ for calibration samples and test samples, we can apply Algorithm 1 or Algorithm 2 to obtain the conformal prediction.
@@
In this setting, the index sets can overlap, meaning that different paths can contain the same edge. However, in a large network without bottlenecks, two shortest paths with random sources and targets share the same edge with a small probability. Therefore, Theorem 2 can guarantee that the coverage of the prediction set is very close to $1-\alpha$.
@@
Road Traffic in Anaheim and Chicago (Bar-Gera, Stabler, and Sall 2023). This dataset is used to predict the traffic flow along certain roads based on the traffic flow along other roads. The predicted traffic flow is then utilized as edge cost for shortest path planning. This scenario corresponds to the situation of subsets with overlaps, where the edges (roads) in different subsets (shortest paths) may overlap with each other. We collect 2000 shortest paths by randomly sampling $(s,t)$ pairs. The Anaheim dataset consists of 413 nodes and 858 edges, and the Chicago dataset consists of 541 nodes and 2150 edges. We adopt a similar procedure from (Jia and Benson 2020; Huang et al. 2024), and allocate 50%, 10%, and 40% for training, validation, and calibration and testing.
7.3 Baseline
@@
This method corresponds to the approach described in Section 3. To predict $\sum_{i\in S^{\textrm{test}}_{k}}y_{i}$, we sample $K+1$ groups of samples from the calibration sets, where each group contains $|S^{\textrm{test}}_{k}|$ samples. Let $S_{1},\dots,S_{K},S_{K+1}$ be the index sets of these groups. The prediction set is then given by (1).
It is important to note that in this method, the indices of $S_{1},\dots,S_{K},S_{K+1}$ may not belong to the same class, which means that this approach may not be appropriate in all cases. However, the method is always applicable, regardless of the class composition of the sampled groups.
@@
Then, the prediction set for $\sum_{i\in S^{\textrm{test}}_{k}}y_{i}$ is given by
$\left[\sum_{i\in S^{\textrm{test}}_{k}}\widehat{y}_{i}+z_{\frac{\alpha}{2}}\sqrt{|S^{\textrm{test}}_{k}|}\widehat{\sigma},\sum_{i\in S^{\textrm{test}}_{k}}\widehat{y}_{i}+z_{1-\frac{\alpha}{2}}\sqrt{|S^{\textrm{test}}_{k}|}\widehat{\sigma}\right],$
@@
The Bonferroni correction has been employed in various conformal methods (Fisch et al. 2021; Jin and Candès 2023; Cleaveland et al. 2024). As mentioned in Section 2.2, it is also applicable to our problem. While the Bonferroni correction is a valid approach, it often lacks efficiency in the context of our problem.
7.4 Results
@@
Figure 2: Results for constructing prediction sets for subsets without overlaps (Section 7.1). CIA achieves guaranteed $1-\alpha$ coverage as well as optimal efficiency on various datasets.
Figure 3: Results for constructing prediction sets for subsets with overlaps (Section 7.2). CIA achieves close to $1-\alpha$ coverage and optimal efficiency on two road traffic datasets.
For these five datasets, the left plots present the desired miscoverage rate $\alpha$ versus the true coverage rate. The closer the curve aligns with the line $1-\alpha$, the easier it is for the method to achieve the desired coverage. Our proposed method, CIA, is very close to the line $1-\alpha$. There is a small gap in the result of the Anaheim dataset. This result aligns with Theorem 2. Compared to the Chicago dataset, Anaheim is a smaller traffic network. The probability of two shortest paths overlapping each other can affect the validity. This issue is much less obvious in a larger traffic network like Chicago. The quantity $\delta$ from the theorem, along with the coverage gap, is provided in the appendix. Recall that the group-sample conformal method is proposed in Section 3. This is also a conformal prediction method, so its validity is also close to the desired line $1-\alpha$ in the results of Figure 2. However, the validity is much less reliable in the results of Figure 3. This is because edge weights from random samples are not exchangeable with the edge weights on the shortest path. The method of Bonferroni correction has good validity in the example of the community dataset. The reason is that under the current setting of calibration and test set ratio, each subset contains one or two samples, which is very close to the case of classical conformal prediction. In other datasets, Bonferroni correction has too high coverage. For normal confidence intervals, since the assumption is not true in general, the coverage is much lower than the desired one. In conclusion, our proposed method has the most reliable validity in all datasets that we have studied.
@@
8 Conclusion
In this paper, we have introduced conformalized interval arithmetic with symmetric calibration, to tackle the challenges of conformal prediction for interdependent data. Our approach offers methodological contributions and theoretical guarantees for achieving the desired coverage under reasonable exchangeability assumptions. Extensive experiments on real-world datasets demonstrate the validity and efficiency of our method, outperforming others even when they fail to ensure valid coverage. We hope our core ideas and techniques will foster the development of more robust conformal prediction methods across various domains.
9 Acknowledgments
The work described in this paper was partially supported by a grant from City University of Hong Kong (Project No.9610639). We thank the reviewers for their insightful suggestions. We also appreciate the helpful discussions with Andro Sabashvili about Theorem 1.
@@ Appendix A Proof of Theorem 1
We first consider a fixed permutation and a fixed partition of calibration and test set. Let $r^{\textrm{cal}}_{k}$ be the rank of $s^{\textrm{cal}}_{k}$ in $s^{\textrm{cal}}_{k}$ for $k\in[K+1]$. We exchange the calibration set and test set in $S_{k}$, i.e., we swap $S^{\textrm{test}}_{K+1}$ and $S^{\textrm{cal}}_{k}$. Under the symmetry assumption on the calibration indices and test indices, this swapping procedure corresponds two possible partitions with the same probability. Let $r^{\textrm{test}}_{K+1}$ be the rank of $s^{\textrm{test}}_{K+1}$ in $\{s^{\textrm{cal}}_{k}:k=1,\dots,K\}\cup\{s^{\textrm{test}}_{K+1}\}$ after swapping. Under the assumption of Theorem 1, the scores of the calibration set remains the same. Hence $r^{\textrm{test}}_{K+1}=r^{\textrm{cal}}_{K+1}$. $r^{\textrm{test}}_{K+1}$ has the same distribution as $r^{\textrm{cal}}_{K+1}$. Hence,
@@
The event $r^{\textrm{test}}_{K+1}\leq\lceil(1+K)(1-\alpha)\rceil$ is equivalent to $\sum_{i\in S^{\textrm{test}}_{K+1}}y_{i}\in\mathcal{C}(S^{\textrm{test}}_{K+1})$ using the definition of $Q_{1-\alpha}$. This completes the proof.
@@ Appendix B Proof of Theorem 2
Appendix C Summary of Methods and Results
C.1 Methods
CIA Variants
  1. 1.
    $s^{\textrm{cal}}_{k}:=\left|\sum_{i\in S^{\textrm{cal}}_{k}}(y_{i}-\widehat{y}_{i})\right|,k\in[K].$ (8)
  2. 2.
    $s^{\textrm{cal}}_{k}=\max\left\{\sum_{i\in S^{\textrm{cal}}_{k}}(\widehat{y}_{i}^{\alpha/2}-y_{i}),\sum_{i\in S^{\textrm{cal}}_{k}}(y_{i}-\widehat{y}^{1-\alpha/2}_{i})\right\},k\in[K].$ (9)
  3. 3.
  4. 4.
Baseline Using Conformal Prediction
  1. 1.
  2. 2.
  3. 3.
  4. 4.
Baseline Using Normal Approximation
  1. 1.
  2. 2.
C.2 Code Implementation
The attached code contains the implementation of CIA and other baseline methods. The main file, main.py, runs experiments on the bike, community, and MEPS datasets, while plot_results.py generates plots of the results. The code relies on dependencies such as Python 3.x, PyTorch, NumPy, Pandas, Matplotlib, and Scikit-learn.
The Split Conformal Method uses a neural network with three fully connected layers, and Conformal Quantile Regression employs quantile regression forests for predictions. For experiments with subsets that have overlaps, the gnn folder contains additional code, including a directed graph autoencoder implemented in digae_traffic.py and autoencoder.py. Required packages include networkx, geopandas, and torch-geometric.
C.3 Results
We present the comparison results of the aforementioned methods across various datasets discussed in the two applications in the experiment section of the main paper. The results demonstrate the performance of each method in terms of coverage probability and set size.
Appendix D Analysis of CIA on Overlapping Subsets